OpenStudy (anonymous):

9^3 * 9^4 plz help !!!

OpenStudy (anonymous):

so what 9x9x9?

OpenStudy (anonymous):

what's*

OpenStudy (anonymous):

you can use a calculator if you want

OpenStudy (anonymous):

so that would be 729

OpenStudy (anonymous):

next you would have to solve 9x9x9x9

OpenStudy (anonymous):

so that would be 6561

OpenStudy (anonymous):

so now we have 729 times 6561

OpenStudy (anonymous):

And that is 4882969

OpenStudy (anonymous):

so that would give you 4,782,969

OpenStudy (anonymous):

there you go :)

OpenStudy (anonymous):

Yea sort my calculations were wrong

OpenStudy (anonymous):

***sorry

OpenStudy (anonymous):

its fine

OpenStudy (anonymous):

So my answer Is 4,782,969?

OpenStudy (anonymous):

Yep!

OpenStudy (anonymous):

Thanks XD :)

OpenStudy (anonymous):

:D

OpenStudy (jango_in_dtown):

9^3*9^4=9^(3+4)=9^7

OpenStudy (anonymous):

It's 3^4

OpenStudy (anonymous):

@jango_IN_DTOWN

OpenStudy (jango_in_dtown):

what is the correct question?

OpenStudy (anonymous):

Idk